OpenStudy (anonymous):

Medal for help! A constant force of F = 47i + 30j moves an object along a vector D = 13i - 8j, where units are in pounds and feet. Find the work done. a. 851 foot-pounds b. 371 foot-pounds c. -851 foot-pounds d. 851.1 foot-pounds

OpenStudy (michele_laino):

we have to compute the scalar product between force and distance

OpenStudy (michele_laino):

hint: \[\Large W = 47 \times 13 + 30 \times \left( { - 8} \right) = ...?\]

OpenStudy (michele_laino):

where W indicates the work done by your force

OpenStudy (anonymous):

Ah I see so W = 371